# Spectral Moment Formulae for $\hbox{GL}(3)\times\hbox{GL}(2)$ L-functions II: The Eisenstein Case

### v2

@@ 3.3.4. Automorphic $L$-functions
Suppose $\Phi$ is given by (3.31) and $\phi$ is an automorphic form of $\Gamma_{2}$.
@@
Recall (3.23) and (2.8). Since $\widetilde{\Phi}^{\mathrm{ND}}$ is left-invariant by $\begin{pmatrix}\gamma&\\ &1\end{pmatrix}$ for any $\gamma\in\Gamma_{2}$, the pairing in (3.37) is well-defined. The rest of the proof follows [Kw23, Proposition 5.15] (or [Gold, Chapter 12.2]). ∎
Definition 3.21.
For $\sigma>3/2$, the standard $L$-function of $\Phi$ is defined by
$L\left(s,\Phi\right)\ :=\ \sum_{n=1}^{\infty}\ \frac{\mathcal{B}_{\Phi}(1,n)}{n^{s}}.$ (3.39)
By [Gold, Theorem 10.8.6], we have
@@
for $\sigma>3/2$, and we have $\{\beta_{\phi,1}(p),\beta_{\phi,2}(p)\}=\{p^{\mu},p^{-\mu}\}$ if $\phi=E^{*}(*;\mu)$. The results follow from Cauchy’s identity in the form [Gold, Proposition 7.4.12], i.e.,
@@ 4. The integral transform
Let $h=H^{\flat}$ and $G_{\alpha}$ as in (3.13) and (3.8). In [Kw25, Section 5.3], we show that the integral transform associated with Theorem 1.3 admits the following simple integral representation:  [^5] The same transform was defined in [Kw23, Kw25] under the notation $(\mathcal{F}_{\Phi}H)(s_{0},\,s)$. In fact, it depends only on the spectral parameters $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$, and hence, we adopt the present notation.
@@
$(\mathcal{F}^{\pm}_{\Phi}H)\left(s_{0},s\right)\,:=\,\int_{0}^{\infty}\int_{0}^{\infty}\,h(y_{1}^{-1})(y_{0}^{2}y_{1}^{-1})^{s-\frac{1}{2}}\int_{0}^{\infty}\,W_{\rho(w_{\ell})\Phi}\left[\begin{pmatrix}\pm X&&\\ 1&1&\\ &&1\end{pmatrix}\mathbf{y}\right]X^{s_{0}-1}\,d^{\times}X\,d^{\times}y_{0}\,dy_{1};$ (4.1)
@@
as in [Kw23, Section 4B]. The innermost integral over $X$ reflects the non-archimedean structure of the dual moment (1.9), as expected. By Lemma 3.4, it follows that $(\mathcal{F}_{\alpha}H)(s_{0},s)$ is equal to the limit of   $\sum_{\pm}\,(\mathcal{F}_{\alpha}^{\pm}H)(s_{0},\,s;\,\phi)$ as $\phi\to\pi/2$-, where
@@
More explicitly, we have  [^6] This corrects a minor typo in [Kw23, eq. (9-13)], where the factor $\pi^{-s_{0}}$ is missing.
@@
See [Kw23, Propositions 8.1 and 9.1].  [^7] The first part of Proposition 4.1 is obtained by inspecting (4).
@@
If one were analyzing $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ with the Kuznetsov formulae, one must treat the “$J$-Bessel” (“same-sign”) and the “$K$-Bessel” (“opposite-sign”) pieces separately (cf. [Mo97, Theorems 2.2 & 2.4]). By contrast, our period integral approach yields a single off-diagonal piece involving the $\hbox{PGL}(3)$ Whittaker function (see (6.32)), and this results in much simpler computations.
Observant readers may wonder about the formulation of Theorem 1.3: we symmetrize the spectral side to obtain a clean description of the main terms, and the dual side should likewise involve the completed $L$-functions. This is indeed the case and is implicit in [Kw23, Theorem 10.6]. Furthermore, this permits us to rewrite the integral transform symmetrically in terms of the Gauss hypergeometric function (2.5). Although not needed later in the article, we believe that this observation is of independent interest. Let us take $s=1/2$ and $\gamma(x):=\Gamma(x)\Gamma(\frac{1}{2}-x)^{-1}$. We first recall:
@@
([Kw23, Theorem 10.6]) Suppose $\operatorname{Re}s_{0}=1/2$. Then
@@
In [Kw23], the proof of Theorem 4.3 invokes [IS07, Proposition 2 & Corollary 4] and [Ish19, Lemma 1.3] on Mellin–Barnes integrals, both of which are elementary consequences of the first Barnes lemma:
@@
The ingredient in [IS07] is a recursion expressing the Whittaker function of $\overline{\mathrm{G}}_{3}$ in terms of that of $\overline{\mathrm{G}}_{2}$, which is subsequently generalized by [Jac09] and [Hum25].
@@
$\left(\mathcal{F}_{\alpha}H\right)(s_{0})\ =\$$2\,\Gamma_{\mathbb{R}}(1-s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(s_{0}+\alpha_{i})\int_{0}^{\infty}\,x^{1/2}\,\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}}{2},\;\frac{1-s_{0}-\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\mathbf{F}\left(\begin{matrix}\frac{s_{0}+\alpha_{2}}{2},\;\frac{s_{0}+\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)$
$\hskip 40.0pt\cdot\,\int_{(0)}\ H(\mu)\prod_{\pm}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(1/2\pm\mu\pm\alpha_{1})\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\,\,d^{W}\mu\ d^{\times}x.$ (4.11)
@@
Substituting the last expression into (4.8) and rewriting with $x\to x^{-1}$ and $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\,\Gamma(s/2)$, our desired result follows. ∎
@@
[…]
@@
Also, we wish to point out several other forms of the integral transforms obtained by Humphries–Khan [HK22+, eq. (3.5)] and Biró [Bi22+, p. 5]. Readers should consult their works for the formulae and the relevant regularity assumptions.
@@ 5. Structures of the main terms
The authors of [CFK+05] proposed an elegant heuristic for obtaining the full sets of main terms for very general classes of moments of $L$-functions. It avoids repeated Taylor/Laurent expansions and proliferation of transcendental constants. More importantly, it encodes the underlying combinatorics and symmetries, and shows remarkable agreement with random matrix theory, specifically, with moments of characteristic polynomials.
@@ 5.1. CFKRS for orthogonal symmetry
For the shifted cubic moment $\mathfrak{M}_{\alpha}^{(3)}(s;H)$, the CFKRS conjecture predicts $2^{3}=8$ main terms in total, which can be classified as follows [CF22+]. Firstly, the diagonal term, or the “0-swap” term, is of the form:
@@
The main terms of $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ are classified by the rightmost expression in (3.24): the sums over
@@
Observe that the “4=3+14=3+1” decomposition adopted in this work fits more naturally with the structures of the main terms for $\mathfrak{M}_{\alpha}^{(3)}(s;H)$ than the conventional “4=2+24=2+2” counterpart (see [Iv02, Fr20, Mo93, Mo97], for instance).
A salient feature in many explicit archimedean computations is the complete reduction of the Mellin–Barnes integrals to ratios of products of $\Gamma$-factors, provided the arguments lie in certain “nice” configurations. This phenomenon typically reflects deeper structures of the integrals. In our case, the reductions of the integral transforms $(\mathcal{F}_{\alpha}H)(s_{0},s)$ and $J_{w,\alpha}(s;h)$ correspond exactly to the CFKRS conjecture for $\mathfrak{M}_{\alpha}^{(3)}(s;H)$. This is the content of Theorem 1.4.
@@
Another example of complete reduction of Mellin–Barnes integrals occurs in [St01, Theorem 3.3]. It corresponds to the exterior-square lifting for $\mathrm{GL}(n)$.
@@ 5.2. CFKRS for unitary symmetry
where $\tau_{A}(m):=\sum_{d_{1}d_{2}=m}\ d_{1}^{-\alpha_{1}}d_{2}^{-\alpha_{2}}$ if $A=\{\alpha_{1},\alpha_{2}\}$, and similarly for $\tau_{B}(n)$ if $B=\{\beta_{1},\,\beta_{2}\}$. The series (5.2) comes up naturally as the recipe of [CFK+05, Sections 1.7 and 2] captures only the diagonal. For $\sigma>0$, we have
@@
and the infinite Euler product $\widetilde{\mathcal{A}}_{AB}(s)$ converges absolutely on $\sigma>-\delta$ for some small $\delta>0$ (see [CFK+05, Theorem 2.4.2.1]). The full set of the main terms for the shifted moment
@@
The formulation (5.5) generalizes to any finite sets $A$ and $B$ of shifts with $|A|=|B|$, where $\tau_{A}$, $\tau_{B}$ are defined similarly, and the upper limit “22” of the $\ell$-sum is replaced by $|A|$ $(=|B|)$. However, the corresponding Euler product $\widetilde{\mathcal{A}}_{AB}(s)$ does not admit a simple closed form in general ([CIS12]).

### v3

@@ 3.3.4. Automorphic $L$-functions
Suppose $\Phi$ is given by (3.33) and $\phi$ is an automorphic form of $\Gamma_{2}$ as in Section 3.3.1. Recall that
Definition 3.19.
For $\sigma>3/2$, the standard $L$-functions of $\phi$ and $\Phi$ are defined, respectively, by
$L(s,\phi)\ :=\ \sum_{n=1}^{\infty}\ \frac{\mathcal{B}_{\phi}(n)}{n^{s}}\hskip 15.0pt\text{and}\hskip 15.0ptL(s,\Phi)\ :=\ \sum_{n=1}^{\infty}\ \frac{\mathcal{B}_{\Phi}(1,n)}{n^{s}}.$ (3.38)
By [31, Theorem 10.8.6], we have
When $\phi$ is an even cusp form, we have the functional equation (see [31, Proposition 3.13.5]):
$\Lambda(s,\phi)\ =\ \prod_{\pm}\,\Gamma_{\mathbb{R}}(s\pm\mu)L(s,\phi)\ =\ \Lambda(1-s,\phi).$ (3.40)
@@
Recall (3.23) and (2.11). Since $\widetilde{\Phi}^{\mathrm{ND}}$ is left-invariant by $\begin{pmatrix}\gamma&\\ &1\end{pmatrix}$ for any $\gamma\in\Gamma_{2}$, the pairing in (3.42) is well-defined. The rest of the proof follows [49, Proposition 5.15] (or [31, Chapter 12.2]). ∎
@@
for $\sigma>3/2$, and we have $\{\beta_{\phi,1}(p),\beta_{\phi,2}(p)\}=\{p^{\mu},p^{-\mu}\}$ if $\phi=E^{*}(*;\mu)$. The results follow from Cauchy’s identity in the form [31, Proposition 7.4.12], i.e.,
@@ 4. The integral transform
Let $h=H^{\flat}$ and $G_{\alpha}$ as in (3.13) and (3.8). In [50, Section 5.3], we show that the integral transform associated with Theorem 1.3 admits the following simple integral representation:  [^6] The same transform was defined in [49, 50] under the notation $(\mathcal{F}_{\Phi}H)(s_{0},\,s)$. In fact, it depends only on the spectral parameters $\alpha=(\alpha_{1},\alpha_{2},\alpha_{3})$, and hence, we adopt the present notation.
@@
$(\mathcal{F}^{\pm}_{\Phi}H)\left(s_{0},s\right):=\int_{0}^{\infty}\int_{0}^{\infty}h(y_{1}^{-1})(y_{0}^{2}y_{1}^{-1})^{s-\frac{1}{2}}\int_{0}^{\infty}W_{\rho(w_{\ell})\Phi}\left[\begin{pmatrix}\pm x&&\\ 1&1&\\ &&1\end{pmatrix}\begin{pmatrix}y_{0}y_{1}&&\\ &y_{1}&\\ &&1\end{pmatrix}\right]x^{s_{0}-1}\,d^{\times}x\,d^{\times}y_{0}\,dy_{1};$ (4.1)
@@
as in [49, Section 4B]. The innermost integral over $x$ reflects the non-archimedean structure of the dual moment (1.7), as expected. By Lemma 3.4, it follows that $(\mathcal{F}_{\alpha}H)(s_{0},s)$ is equal to the limit of   $\sum_{\pm}\,(\mathcal{F}_{\alpha}^{\pm}H)(s_{0},\,s;\,\phi)$ as $\phi\to\pi/2$-, where
@@
More explicitly, we have  [^7] This corrects a minor typo in [49, eq. (9-13)], where the factor $\pi^{-s_{0}}$ is missing.
@@
See [49, Propositions 8.1 and 9.1].  [^8] The first part of Proposition 4.1 is obtained by inspecting (4).
@@
If one were analyzing the moment $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ with the Kuznetsov formulae, one must treat the “$J$-Bessel” (“same-sign”) and the “$K$-Bessel” (“opposite-sign”) pieces separately (cf. [60, Theorems 2.2 & 2.4]). By contrast, our period integral approach yields a single off-diagonal piece involving the $\hbox{PGL}(3)$ Whittaker function (see (6.34)), and this results in much simpler computations.
Observant readers may wonder about the formulation of Theorem 1.3. Since we symmetrize the spectral side to obtain a clean description of the main terms, one may likewise expect the dual side to involve the completed $L$-functions. This is indeed the case. It is implicit in [49, Theorem 10.6], and is important for reasons to be explained in Remark 4.5. Furthermore, this permits us to rewrite the integral transform symmetrically in terms of the Gauss hypergeometric function (2.5). Although not needed later in the article, we believe that this observation is of independent interest.
Let us take $s=1/2$ and $\gamma(x):=\Gamma(x)\Gamma(1/2-x)^{-1}$. We first recall:
@@
([49, Theorem 10.6]) Suppose $\operatorname{Re}s_{0}=1/2$. Then
@@
In [49], the proof of Theorem 4.3 invokes [40, Proposition 2 & Corollary 4] and [41, Lemma 1.3] on Mellin–Barnes integrals, both of which are elementary consequences of the first Barnes lemma:
@@
The ingredient in [40] is a recursion expressing the Whittaker function of $\overline{\mathrm{G}}_{3}$ in terms of that of $\overline{\mathrm{G}}_{2}$, which is subsequently generalized by [44] and [37].
@@
$\left(\mathcal{F}_{\alpha}H\right)(s_{0})\ =\$$4\,\Gamma_{\mathbb{R}}(1-s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(s_{0}+\alpha_{i})\int_{0}^{\infty}\,\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}}{2},\;\frac{1-s_{0}-\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)\mathbf{F}\left(\begin{matrix}\frac{s_{0}+\alpha_{2}}{2},\;\frac{s_{0}+\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)$
$\hskip 40.0pt\cdot\,\int_{(0)}\ H(\mu)\prod_{\pm}\,\prod_{\pm}\,\Gamma_{\mathbb{R}}(1/2\pm\mu\pm\alpha_{1})\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{2}\right)\,\,d^{W}\mu\ dx.$ (4.11)
@@
$\mathcal{K}_{\alpha}(s_{0};\,\mu)\,=\,\frac{1}{\sqrt{\pi}}$ $\,\prod_{\pm}\,\Gamma\Big(\frac{1}{4}+\frac{\alpha_{1}\pm\mu}{2}\Big)\int_{0}^{\infty}\,x^{-\frac{1}{2}-\frac{\alpha_{1}}{4}}\,\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x^{-1}\right)$
@@
Substituting the last expression into (4.8) and rewriting with $x\to x^{-2}$ and $\Gamma_{\mathbb{R}}(s):=\pi^{-s/2}\,\Gamma(s/2)$, our desired result follows. ∎
Remark 4.5.
Here we check that the integral transform (4.4) has the expected symmetry. Recall our main result (1.7) with $s=1/2$. It is clear that the sum $(\mathcal{R}_{\alpha}+\mathcal{R}_{-\alpha})(1/2;H)$ is invariant under $\alpha\to-\alpha$. The same is true for the terms  [^10] In fact, they enjoy a larger set of invariance $\alpha\to(\epsilon_{1}\alpha_{1},\epsilon_{2}\alpha_{2},\epsilon_{3}\alpha_{3})$ with $\epsilon_{i}=\pm$, and the significance will be discussed in Section 5.1. $\sum_{i=1}^{3}\,\mathcal{M}_{-\alpha}^{i}(1/2;H)$ and $\mathfrak{M}_{-\alpha}^{(3)}(1/2;H)$; for the latter, this follows from the functional equations (2.3) and (3.40).
Now, applying (2.3) to each of the four copies of the $\zeta$-function in the dual moment of (1.7), the integral transform should satisfy the functional equation
$\frac{(\mathcal{F}_{\alpha}H)(s_{0})}{(\mathcal{F}_{-\alpha}H)(1-s_{0})}\,=\,\frac{\Gamma_{\mathbb{R}}(1-s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(s_{0}+\alpha_{i})}{\Gamma_{\mathbb{R}}(s_{0})\prod_{i=1}^{3}\,\Gamma_{\mathbb{R}}(1-s_{0}-\alpha_{i})}.$ (4.12)
We prove (4.12) directly as follows. Recall the Euler transformation:
$\mathbf{F}\left(\begin{matrix}a,\;b\\ c\end{matrix}\Bigm|z\right)\,=\,(1-z)^{c-a-b}\,\mathbf{F}\left(\begin{matrix}c-a,\;c-b\\ c\end{matrix}\Bigm|z\right).$ (4.13)
It follows that
$\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}}{2},\;\frac{1-s_{0}-\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\,$ $=\,(1+x)^{s_{0}+\frac{\alpha_{1}}{2}-\frac{1}{2}}\,\mathbf{F}\left(\begin{matrix}\frac{s_{0}}{2},\;\frac{s_{0}+\alpha_{1}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right),$
$\mathbf{F}\left(\begin{matrix}\frac{s_{0}+\alpha_{2}}{2},\;\frac{s_{0}+\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\,$ $=\,(1+x)^{\frac{1}{2}-s_{0}+\frac{\alpha_{1}}{2}}\mathbf{F}\left(\begin{matrix}\frac{1-s_{0}-\alpha_{2}}{2},\;\frac{1-s_{0}-\alpha_{3}}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right),$
$\mathbf{F}\left(\begin{matrix}\frac{1}{4}+\frac{\alpha_{1}+\mu}{2},\;\frac{1}{4}+\frac{\alpha_{1}-\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right)\,$ $=\,(1+x)^{-\alpha_{1}}\,\mathbf{F}\left(\begin{matrix}\frac{1}{4}-\frac{\alpha_{1}}{2}-\frac{\mu}{2},\;\frac{1}{4}-\frac{\alpha_{1}}{2}+\frac{\mu}{2}\\ \frac{1}{2}\end{matrix}\Bigm|-x\right),$
where we used $\alpha_{1}+\alpha_{2}+\alpha_{3}=0$ several times. The result follows from the observation that
$(1+x)^{s_{0}+\frac{\alpha_{1}}{2}-\frac{1}{2}}(1+x)^{\frac{1}{2}-s_{0}+\frac{\alpha_{1}}{2}}(1+x)^{-\alpha_{1}}\,=\,1.$
@@
${\mathcal{W}(\widetilde{\Pi})}$$\mathcal{W}(\Pi)$$\big\{H(\pi):\ \widehat{[\overline{\mathrm{G}}_{2}]}\ \rightarrow\ \mathbb{C}\big\}$$\big\{h(z):\ \widehat{\mathbb{R}}_{\mathrm{unit}}^{\times}\ \rightarrow\ \mathbb{C}\big\}$$\scriptstyle{\mathcal{F}_{\alpha}}$
@@
Also, we wish to point out several other forms of the integral transforms obtained by Humphries–Khan [36, eq. (3.5)] and Biró [5, p. 5]. Readers should consult their works for the formulae and the relevant regularity assumptions.
@@ 5. Structures of the main terms
The authors of [22] proposed an elegant heuristic for obtaining the full sets of main terms for very general classes of moments of $L$-functions. It avoids repeated Taylor/Laurent expansions and proliferation of transcendental constants. More importantly, it encodes the underlying combinatorics and symmetries, and shows remarkable agreement with random matrix theory, specifically, with moments of characteristic polynomials.
@@ 5.1. CFKRS for orthogonal symmetry
For the shifted cubic moment $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$, the CFKRS conjecture predicts $2^{3}=8$ main terms in total, which can be classified as follows [18]. Firstly, the diagonal term, or the “0-swap” term, is of the form:
@@
The main terms of $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ are classified by the rightmost expression in (3.24): the sums over
@@
Observe that the “4=3+14=3+1” decomposition adopted in this work fits more naturally with the structures of the main terms for $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$ than the conventional “4=2+24=2+2” counterpart (see [43, 30, 59, 60], for instance).
A salient feature in many explicit archimedean computations is the complete reduction of the Mellin–Barnes integrals to ratios of products of $\Gamma$-factors, provided the arguments lie in certain “nice” configurations. This phenomenon typically reflects deeper structures of the integrals. In our case, the reductions of the integral transforms $(\mathcal{F}_{\alpha}H)(s_{0},s)$ and $J_{w,\alpha}(s;h)$ correspond exactly to the CFKRS conjecture for $\mathfrak{M}_{-\alpha}^{(3)}(s;H)$. This is the content of Theorem 1.4.
@@
Another example of complete reduction of Mellin–Barnes integrals occurs in [70, Theorem 3.3]. It corresponds to the exterior-square lifting for $\mathrm{GL}(n)$.
@@ 5.2. CFKRS for unitary symmetry
where $\tau_{A}(m):=\sum_{d_{1}d_{2}=m}\ d_{1}^{-\alpha_{1}}d_{2}^{-\alpha_{2}}$ if $A=\{\alpha_{1},\alpha_{2}\}$, and similarly for $\tau_{B}(n)$ if $B=\{\beta_{1},\,\beta_{2}\}$. The series (5.2) comes up naturally as the recipe of [22, Sections 1.7 and 2] captures only the diagonal. For $\sigma>0$, we have
@@
and the infinite Euler product $\widetilde{\mathcal{A}}_{AB}(s)$ converges absolutely on $\sigma>-\delta$ for some small $\delta>0$ (see [22, Theorem 2.4.2.1]). The full set of the main terms for the shifted moment
@@
The formulation (5.5) generalizes to any finite sets AA and $B$ of shifts with $|A|=|B|$, where $\tau_{A}$, $\tau_{B}$ are defined similarly, and the upper limit “22” of the $\ell$-sum is replaced by $|A|$ $(=|B|)$. However, the corresponding Euler product $\widetilde{\mathcal{A}}_{AB}(s)$ does not admit a simple closed form in general ([23]).